OpenStudy (anonymous):

3 2/9 times 6/7 as a mixed number in simplest form

OpenStudy (jdoe0001):

\(\bf 3\cfrac{2}{9}\cdot \cfrac{6}{7}\implies \cfrac{3\cdot 9+2}{9}\cdot \cfrac{6}{7}\implies \cfrac{(3\cdot 9+2)\cdot 6}{9\cdot 7}\)

OpenStudy (anonymous):



OpenStudy (jdoe0001):

ahemm you kinda bypassed PEMDAS ^ parentheses first

OpenStudy (jdoe0001):

\(\bf 3\cfrac{2}{9}\cdot \cfrac{6}{7}\implies \cfrac{3\cdot 9+2}{9}\cdot \cfrac{6}{7}\implies \cfrac{({\color{red}{ 3\cdot 9+2}})\cdot 6}{9\cdot 7}\implies \cfrac{{\color{red}{ 29}}\cdot 6}{63}\)

OpenStudy (anonymous):